OpenStudy (anonymous):

How would i find cos(225) using sum and difference identities?

OpenStudy (anonymous):

since \(225=180+45\) you could use that one

OpenStudy (anonymous):

\[\cos(225)=\cos(180+45)=\cos(180)\cos(45)-\sin(180)\sin(45)\] is a start

OpenStudy (anonymous):

180 is an obtuse angle, dont I have to use acute angles?

OpenStudy (anonymous):

no it makes no difference at all

OpenStudy (anonymous):

doubt you are going to find two acute angles that add up to \(225\) in any case

OpenStudy (anonymous):

ok thank you (:

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

so would my answer be - square root of 2 over 2